OpenStudy (anonymous):

Find the slope between points (-4, 8) and (-9, 0).

OpenStudy (anonymous):

y-y=m(x-x)

OpenStudy (hoblos):

slope between (x1,y1) and (x2,y2) is (y2-y1)/(x2-x1)

OpenStudy (anonymous):

8+9=m(-4-0). : 17=-4m........17/-4=m.

OpenStudy (anonymous):

the choices are : A -8/5 B 4/9 C 8/5 D undefined slope

OpenStudy (hoblos):

m= (0-8)/(-9+4) = (-8)/(-5) = 8/5

OpenStudy (anonymous):

thank you!

OpenStudy (hoblos):

my pleasure XD